# Zhonghong Kuang Renmin U Yi Liu

### v1

@@ 3.2 Characterization
[…]
@@
Given a contract $(\bm{s},\bm{p})$, from the standpoint of any period $k$, sticking to the contract gives the principal roughly
@@
[…]
@@ 3.3 Pareto Efficiency
Expert’s time-00 value ($W_{0}$)
Principal’s time-00 profit ($\Pi_{0}$)
[…]
Pareto frontier
Figure 3: Pareto Frontier
@@
Our notion of Pareto efficiency considers only the principal and the expert. Nevertheless, Theorem 2 holds verbatim even if the novice’s welfare is included in the definition. This is because the novice’s stage payoff, $v(\cdot)$, is an increasing function (Assumption 1) and $\bm{s}^{*}$ dominates any other implementable $\bm{s}$ (Lemma 6).
3.4 Application to the Chinese Auto Industry
@@
By the 1970s, the quid pro quo policy, which required multinational companies to transfer technology in return for market access, had become a common practice in many developing countries (Holmes, McGrattan and Prescott, 2015). After launching its economic reform in 1978, China urgently needed to learn advanced technologies and explore domestic markets, especially in the auto industry. Since then, China adopted the quid pro quo policy and followed it through until very recently. The requirement for technology transfer was explicit before 2001 (the year China joined the WTO) and became implicit afterwards. Bai, Barwick, Cao and Li (2025) provide empirical evidence for the effectiveness of this policy and discuss the institutional background in detail.
@@
The primary way to implement the quid pro quo policy was through joint ventures (JVs) between international enterprises and domestic partners. The first JV for automobile manufacturing was formed in 1983 between American Motors Corporation and Beijing Jeep Corporation. Other notable examples include the JV between Shanghai Automobile Industry Corporation (SAIC) and Volkswagen established in 1984, and that between SAIC and General Motors (GM) established in 1997. These JVs required foreign automakers to offer know-how and product lines as equity (capped at 50%), while domestic partners provided manufacturing facilities and labor. Consistent with Lemma 5, this particular arrangement helped facilitate substantial knowledge transfer in the initial stage of the partnership.
@@
Establishing JVs was only the first step, as technology transfer was not an overnight process; instead, it typically lasted for years following a foreign firm’s market entry. Take SAIC as an example. Despite its JVs beginning in the 80s and 90s, SAIC lacked the capability of manufacturing passenger vehicles on its own until 2006, when it launched the inaugural independent automotive brand, Roewe; even after that, cars sold under this local brand were generally perceived as lower-end and of inferior quality compared to those under a JV brand. Meanwhile, since these JVs were formed, both central and local governments have continually supported participating foreign companies through a variety of subsidies, including guaranteed government procurements, land price discounts, and tax incentives. [^15] For example, the contract between SAIC and Volkswagen specified that the Shanghai government guaranteed the purchase of Santana vehicles for taxis (Yuan and Broggi, 2025). Similarly, between 2004 and 2015, most taxis in Beijing (around 80,000 vehicles) were produced by Beijing-Hyundai, a joint venture established in 2002. Since 1986, local governments in China have had the authority to lower the costs for land use. Even recently, SAIC-Volkswagen acquired a new land parcel of 95,394 square meters at an ultra-low rate of 1,020 RMB per square meter from the Shanghai Government in 2022. See Article 4 of Provisions of the State Council for the Encouragement of Foreign Investment (English version: http://www.law-lib.com/lawhtm/1986/3910.htm). From 1991 to 2007, eligible foreign enterprises were exempted from income tax during their first two profitable years. Subsequently, they received a 50% reduction in income tax from the third to the fifth year. See Article 8 of The Income Tax Law of the People’s Republic of China for Enterprises with Foreign Investment and Foreign Enterprises (English version: http://www.law-lib.com/lawhtm/1991/7536.htm).
@@
We also observe that some, but not all, international companies have expressed reservations about a complete technology transfer. For example, it was reported that “GM is holding tight to its more valuable technology: Beijing is eager to tap into foreign auto companies’ clean-energy technologies, [b]ut GM doesn’t want to share all its research with its Chinese partner.” [^16] See Sharon Terlep, “Balancing the Give and Take in GM’s Chinese Partnership.” Wall Street Journal, August 19, 2012. In contrast, Volkswagen has appeared more open to collaborations on even its most advanced technologies, working closely with SAIC on developing and deploying innovations regarding electrification and autonomous driving. [^17] See Aschhoff and Ludewig, “All set for future mobility: Volkswagen Group and SAIC Strengthen Long-Standing Partnership with New Joint Venture Agreement,” Volkswagen Group Press Release, November 27, 2024. This divergence may be attributable to the fact that the economic relationship between China and the EU is often considered more stable and less politically sensitive than that between China and the US. As a result, the SAIC-Volkswagen joint venture is arguably associated with a higher discount factor and, as predicted by Proposition 1, is able to sustain a greater level of technology transfer in the long run.
@@ 4 Expert’s Retirement
A crucial feature of the optimal contract in Section 3 is that knowledge transfer does not end in finite time, and when the players are relatively impatient, knowledge cannot be fully transferred even in the long run. This is somewhat undesirable due to the inefficiencies it causes; moreover, in some applications such as employment relationships, having an expert that is as long-lived as the principal (firm) might not be the most reasonable setting to begin with.
@@ 4.1 Analysis
In apprenticeships, it is common for a master to teach his student without reservation when he is close to his retirement, even if the student is a potential competitor. Take independent watchmakers as an example: Under the voluntary guidance of Philippe Dufour and two founders of Greubel Forsey, Michel Boulanger was trained with traditional manual skills and craftsmanship starting in 2007. Later, he put into practice the techniques he had learned by creating a timepiece by hand, which was auctioned for almost 1.5M USD to finance the subsequent transmission of craftsmanship to the next generation. [^19] For more details, see https://timeaeon.org/projects/naissance-dune-montre.
@@
$p_{t}^{*}$ $=\frac{\left(1-\delta^{K-t}\right)[w(s_{t-1})-w(s_{t})]}{1-\delta},\text{ for }t\in\{1,...,K-1\}.$
@@
$\underbrace{\frac{(\delta-\delta^{K-t})[\pi(s_{t+1})-\pi(s_{t})]}{1-\delta}}_{\text{Benefit of $x_{t}$ through stage payoffs}}+\underbrace{\delta^{K-1-t}[C(s_{t})-C(s_{t+1})]}_{\text{Benefit of $x_{t}$ through cost savings}}=\underbrace{\frac{(1-\delta^{K-t})[w(s_{t-1})-w(s_{t})]}{1-\delta}}_{\text{Compensation to the expert}}.\vspace{0.25cm}$ (R-BE)
Like before, the break-even condition requires that, in each period, the principal’s total benefit from the expert’s training must equal the payment she is supposed to make in that period. In this case, the expert’s training in period $t$ not only affects stage payoffs from period $t+1$ through $K-1$, but also the principal’s catch-up cost after the current expert retires.
Then, we can write the principal’s time-00 profit as
@@
[…]
Similar to Lemma 5, the optimal knowledge sequence maximizes training in the very first period (i.e., $x_{0}$) among those that make the players’ incentive constraints binding. Intuitively, the principal’s benefit from all the training starting $t=1$ is fully offset by her payment obligations, but the training at $t=0$ is provided “for free” (i.e., $p_{0}=0$). Thus, profit maximization boils down to maximizing the principal’s rent that results from such free training.
@@
There exists a unique optimal contract with retirement $(\bm{s}^{*},\ \bm{p}^{*})$, and it has the following structure:
@@ 4.2 Effect of Catch-Up Cost
Absent the catch-up cost, knowledge transfer is not possible except in an expert’s final period of the game. This is because the principal would have no incentives to pay in that final period, but then the expert would not want to train the novice in the period before, leading to unraveling. In contrast, with the catch-up cost, if the principal does not pay a retiring expert, the latter can still punish the principal by withholding training in his final period. This credible punishment makes higher payments and better training sustainable from the beginning of the game. The next result formalizes this intuition.
@@
Fix any two catch-up cost functions $C(\cdot)$ and $\hat{C}(\cdot)$ such that $\hat{C}(s)>C(s)$ for $s\in[0,1)$.

### v2

@@ 3.2 Characterization
There exists a unique optimal contract $(\bm{s}^{*},\ \bm{p}^{*})$. If $\bar{s}^{*}=0$, $(\bm{s}^{*},\ \bm{p}^{*})$ is the trivial contract with no knowledge transfer. If $\bar{s}^{*}\in(0,\ 1]$, $(\bm{s}^{*},\ \bm{p}^{*})$ has the following structure:
@@
Given a contract $(\bm{s},\bm{p})$, in any period $k$, sticking to the contract gives the principal roughly
@@
where $\frac{w(s_{k-1})-w(\bar{s})}{1-\delta}$ is the approximate total payment from period $k$ onward until the knowledge level stabilizes, and $\frac{\delta\pi(\bar{s})}{1-\delta}$ is the approximate value of future knowledge following the principal’s payment at $k$. On the other hand, if the principal were to default, he would only obtain $\frac{\delta\pi(s_{k})}{1-\delta}$ starting k+1k+1.111111Note that the principal can get $\pi(s_{k})$ at $k$ regardless of his payment decision in the current period, so it drops out of the payoff comparison. As a result, increasing $\delta\pi(\bar{s})+w(\bar{s})$ would always help enhance the long-run value of the contract to the principal, making higher payments and hence greater knowledge transfers sustainable.
@@ 3.3 Pareto Efficiency
Expert’s time-0 value ($W_{0}$)
Principal’s time-0 profit ($\Pi_{0}$)
[…]
Pareto frontier
Figure 3: Pareto Frontier
@@
Our notion of Pareto efficiency considers only the principal and the expert. Theorem 2 holds verbatim if the novice’s welfare is also included in the definition, because the novice’s stage payoff, $v(\cdot)$, is an increasing function (Assumption 1) and $\bm{s}^{*}$ uniformly dominates any other implementable $\bm{s}$ (Lemma 6).
4 Applications
In this section, we apply our prior analysis to two cases of large-scale cross-border technology transfers, one in the Chinese auto industry and the other pertaining to the development of Korea Train Express.
4.1 The Chinese Auto Industry
@@
Beginning in the 1970s, many developing countries implemented the quid pro quo policy under which multinational corporations were required to transfer technology as a condition for market access (Holmes, McGrattan and Prescott, 2015). After launching its economic reform in 1978, China urgently needed to learn advanced technologies and explore domestic markets. Shortly after, China adopted the quid pro quo policy in the automobile industry and followed it through until very recently. The requirement for technology transfer was explicit before 2001 (the year China joined the WTO) and became implicit afterwards. Bai, Barwick, Cao and Li (2025) provide empirical evidence for the effectiveness of this policy and discuss the institutional background in detail.
@@
The threat to exclude induced many automakers to participate in a technology transfer process, and the primary way to implement it was through joint ventures (JVs) between international automakers and domestic partners. [^12] In the model, we take participation as given without explicitly considering this pre-game exclusion threat. The first JV for automobile manufacturing was formed in 1983 between American Motors Corporation and Beijing Jeep Corporation. Other notable examples include the JV between Shanghai Automobile Industry Corporation (SAIC) and Volkswagen established in 1984, and that between SAIC and General Motors (GM) established in 1997. The “Automotive Industry Development Policy” issued by China’s State Council mandated that cars manufactured by the JVs should reach the international technological level of the same period. [^13] See Article 16 of Chapter 4 of the Automotive Industry Development Policy. To meet this requirement, foreign automakers offered know-how and product lines as equity (capped at 50%) soon after the policy was implemented, while domestic partners provided manufacturing facilities and labor. Consistent with Lemma 5, this policy helped facilitate substantial knowledge transfer in the initial stage of the partnership.
@@
Establishing JVs was only the first step, as technology transfer was not an overnight process. Instead, it lasted for years via various channels, such as continued training of local workers and management, transfer and joint ownership of patents, and setting up R&D offices within the JVs. Meanwhile, since these JVs were established, both central and local governments have continually supported participating foreign companies through a variety of subsidies, including guaranteed government procurements, land price discounts, and tax incentives. [^14] For example, the contract between SAIC and Volkswagen specified that the Shanghai government guaranteed the purchase of Santana vehicles for taxis (Jia-Zheng and Broggi, 2025). Similarly, between 2004 and 2015, most taxis in Beijing (around 80,000 vehicles) were produced by Beijing-Hyundai, a joint venture established in 2002. Since 1986, local governments in China have had the authority to lower the costs for land use. Even recently, SAIC-Volkswagen acquired a new land parcel of 95,394 square meters at an ultra-low rate of 1,020 RMB per square meter from the Shanghai Government in 2022. See Article 4 of the Provisions of the State Council for the Encouragement of Foreign Investment (English version: http://www.law-lib.com/lawhtm/1986/3910.htm). From 1991 to 2007, eligible foreign enterprises were exempted from income tax during their first two profitable years. Subsequently, they received a 50% reduction in income tax from the third to the fifth year. See Article 8 of the Income Tax Law of the People’s Republic of China for Enterprises with Foreign Investment and Foreign Enterprises (English version: http://www.law-lib.com/lawhtm/1991/7536.htm). In line with our predictions, the observed technology transfer process featured gradual, subsequent transfers after the initial push by the government; in turn, the government also provided continued compensations to international automakers to keep the scheme running.
However, even in the long run, the technological capacity of domestic manufacturers often remained strictly below that of their foreign partners. Take SAIC as an example. Despite its JVs beginning in the 80s and 90s, SAIC lacked the capability of manufacturing passenger vehicles on its own until 2006, when it launched the inaugural independent automotive brand, Roewe. Even after that, cars sold under this local brand were generally perceived as lower-end and of inferior quality compared to those under a JV brand.
@@
More interestingly, different JVs exhibited distinct levels of collaboration, even after holding fixed the same domestic manufacturer. For example, it was reported that “GM is holding tight to its more valuable technology: Beijing is eager to tap into foreign auto companies’ clean-energy technologies, [b]ut GM doesn’t want to share all its research with its Chinese partner.” [^15] See Sharon Terlep, “Balancing the Give and Take in GM’s Chinese Partnership.” Wall Street Journal, August 19, 2012. In contrast, Volkswagen has appeared more open to collaborations even on its most advanced technologies, working closely with SAIC on developing and deploying innovations regarding electrification and autonomous driving. [^16] See Aschhoff and Ludewig, “All set for future mobility: Volkswagen Group and SAIC Strengthen Long-Standing Partnership with New Joint Venture Agreement,” Volkswagen Group Press Releases, November 27, 2024. This divergence may be attributable to the fact that the economic relationship between China and the EU is often considered more stable and less politically sensitive than that between China and the US. As a result, the SAIC-Volkswagen joint venture is arguably associated with a higher discount factor and, as predicted by Proposition 1, is able to sustain a greater level of technology transfer in the long run.
4.2 The Korea Train Express (KTX)
In the 1990s, the government of South Korea (principal) decided to build a high-speed rail (HSR) linking Seoul and Pusan, two of Korea’s largest cities. To avoid continued dependence on foreign technologies and its potential jeopardy to operation and profitability, the Korean government specified technology transfer from foreign contractors (expert) to local companies (novice) as part of its request for proposal. In 1994, the Franco-British consortium Alstom won the contract after defeating Siemens from Germany and Mitsubishi from Japan, the latter of which was eliminated reportedly because it was unwilling to agree to the technology-transfer clause. For detailed institutional background, see Kang et al. (2014).
The technology transfer process between Alstom and KTX resembles our characterization in Section 3. First, the contract was structured around strict phasing, featuring gradual technology transfer and gradual payments (Rouach and Saperstein, 2004). In the first phase, Alstom manufactured 12 trainsets in Belfort, France, and delivered them directly. In the second phase, Alstom supplied components, and Korean companies assembled 34 trainsets domestically. In the third phase, Korean companies achieved full localization with Alstom’s assistance. Meanwhile, payments from the Korean government were also made in a phased manner: Subsequent large-scale procurement funds would only be disbursed after verifying the technology transfer required in each phase, with the government directly testing the trainsets built by local manufacturers.
The Korean government and Alstom also ensured that a substantial amount of technology transfer took place in the initial phase. According to Alstom’s press release, “the technology transfer began with transferring and updating approximately 350,000 high speed train documents: drawings, specifications, manufacturing documents, procedures, purchasing documents, and training documents. Then came the training of Korean engineers to use them: more than 1,000 Koreans were trained in France; over 400 French engineers provided assistance to the Korean production.” [^17] See “Official opening of the first high speed railway line in Korea,” Alstom Press Releases, March 30, 2004.
In this case, Alstom agreed to transfer all existing knowledge to Korea while simultaneously continuing to develop new expertise. On one hand, it agreed to fully transfer 100% of the relevant technologies and permit 95% of production activities to be carried out locally in Korea. This arrangement enabled Korean firms to acquire critical know-how and hands-on experience, laying the groundwork for them to independently manufacture their own trains in the future. On the other hand, to safeguard its technological advantage, Alstom remained dedicated to conducting efficient research and development initiatives and sustaining the rapid advancement of its core technologies (Rouach and Saperstein, 2004).
It is worth noting that South Korea has officially become an exporter of high-speed rail, competing directly with Alstom in the international market. In June 2024, South Korea and Uzbekistan signed a deal to apply KTX technology in Uzbekistan, with the first batch of trains delivered in December 2025.181818See Nam Hyun-woo, “Hyundai Rotem delivers trains for Uzbekistan ahead of schedule,” The Korea Times, December 10, 2025. This marks the first time KTX technology has been exported, stressing the legit concerns of an expert about the intensified competition that arises from knowledge transfers.
@@ 5 Expert’s Retirement
A crucial feature of the optimal contract in Section 3 is that the knowledge transfer process does not end in finite time, and when the players are relatively impatient, knowledge cannot be fully transferred even in the long run. This is somewhat undesirable due to the inefficiencies it causes; moreover, in some applications such as employment relationships, having an expert that is as long-lived as the principal (firm) might not be the most reasonable setting to begin with.
@@ 5.1 Analysis
In apprenticeships, it is common for a master to teach his student without reservation when he is close to his retirement, even if the student is a potential competitor. Take independent watchmakers as an example: Under the voluntary guidance of Philippe Dufour and two founders of Greubel Forsey, Michel Boulanger was trained with traditional manual skills and craftsmanship starting in 2007. Later, he put into practice the techniques he had learned by creating a timepiece by hand, which was auctioned for almost \$1.5 million USD to finance the subsequent transmission of craftsmanship to the next generation. [^20] For more details, see https://timeaeon.org/projects/naissance-dune-montre.
@@
$p_{t}^{*}$ $=\frac{\left(1-\delta^{K-t}\right)[w(s_{t-1})-w(s_{t})]}{1-\delta},\text{ for }t\in\{1,.,K-1\}.$
@@
[…]
Like before, the break-even condition requires that, in each period, the principal’s total benefit from the expert’s training must equal the payment she is supposed to make in that period. In this case, the expert’s training in period $t$ not only affects stage payoffs from period $t+1$ through K1K-1, but also the principal’s catch-up cost after the current expert retires.
Then, we can write the principal’s time-0 profit as
@@
$(\bm{s},\bm{p})$ is an optimal contract with retirement if and only if $\bm{p}$ is given by (R-P*) and $\bm{s}$ solves
@@
Similar to Lemma 5, the optimal knowledge sequence maximizes training in the very first period (i.e., $x_{0}$) among those that make the players’ incentive constraints binding. Intuitively, the principal’s benefit from all the training starting $t=1$ is fully offset by her payment obligations, but the training at $t=0$ is provided “for free” (i.e., $p_{0}=0$). Thus, profit maximization boils down to maximizing the principal’s rent that results from the initial training.
@@
There is a unique optimal contract with retirement $(\bm{s}^{*},\ \bm{p}^{*})$, and it is structured as follows:
@@ 5.2 Effect of Catch-Up Cost
Absent the catch-up cost, knowledge transfer is not possible except in an expert’s final period of the game. This is because the principal would have no incentives to pay in that final period, but then the expert would not want to train the novice in the period before, leading to unraveling. In contrast, with the catch-up cost, if the principal does not pay a retiring expert, the latter can still punish the principal by withholding training in his final period, thereby forcing the catch-up cost onto the principal. This credible punishment makes higher payments and better training sustainable from the beginning of the game. The next result formalizes this intuition.
@@
[…]
@@
5.3 Application to Craftsmanship Preservation by Charitable Foundations
Many traditional crafts are at risk of disappearing because masters (expert) are often reluctant to pass their skills on to outsiders, preferring instead to teach only family members. This preference arises because teaching a family member mitigates the risk of a novice undermining the expert’s future earnings: knowledge transfers to a family member may continue to generate benefits for the family, while an outsider has no obligation to support the retired expert. However, as noted by the United Nations Educational Scientific and Cultural Organization (UNESCO), “if family members or community members are not interested in learning it, the knowledge may disappear because sharing it with strangers violates tradition.”
To protect these intangible cultural heritages, many grants and foundations (principal) provide monetary incentives to craftsmen to pass on knowledge and skills associated with traditional artisanries to future generations outside of individual families. For example, the Michelangelo Foundation provides opportunities for the younger generation of artisans to collaborate with experienced master craftsmen, engaging in creation and learning together. This pairing ensures that these skills can be passed down from the “older generation” to the “new generation.” UNESCO also raised more than \$4 million USD to support such intergenerational transfers. [^21] See “International Assistance from the Intangible Cultural Heritage Fund” (project ID is 199IAS4142).
Similar to our characterization, these programs usually facilitate year-long learning of craftsmanship with substantial payments made to the master craftsmen in a gradual manner. For example, the Traditional Arts Apprenticeship Grant of the Montana Arts will pay \$3,000 directly to the master, and the payment will be in multiple installments throughout the yearly program. The Homo Faber Fellowship, which was jointly developed by the Michelangelo Foundation and Jaeger-LeCoultre (a Swiss luxury watch manufacturer), also provides continuous payments to the master during the 8-month program.
@@ 6 Discussion
6.1 Comparison with Garicano and Rayo (2017)
In Garicano and Rayo (2017) (hereafter “GR”), the expert employs the novice to produce, but the cash-constrained novice initially possesses neither the productive knowledge nor the financial resources to acquire such knowledge. Because of that, the expert is intrinsically motivated to train the novice, but at the same time, he is concerned that the novice may stop producing for him and pursue an outside option whose value increases with the novice’s knowledge level. GR show that the expert-optimal contract is a multi-period apprenticeship, in which the novice is trained gradually over time and receives full knowledge after a finite number of periods. In the initial period, the expert offers a “knowledge gift,” despite the novice’s inability to produce; after that, the novice works without pay until he fully acquires the remaining knowledge; once the apprenticeship concludes, the novice starts to retain all the output from his production. In what follows, we highlight the key differences between our results and those in GR.
Finite vs. Perpetual Transfer Process
In GR, knowledge is fully transferred after a finite number of periods. This outcome arises because, as the amount of remaining knowledge diminishes, a threshold is eventually reached where the novice’s output in a single period exceeds the value of the remaining knowledge. At that point, the expert has an incentive to transfer all remaining knowledge at once to accelerate revenue collection. This result also hinges on the expert’s one-period commitment: If the expert were able to freely choose how much knowledge to transfer after the novice had produced, he would never transfer all remaining knowledge, knowing that the novice would then appropriate the entire surplus. In contrast, in our model with no commitment, the knowledge transfer process proceeds perpetually, with the novice’s knowledge level converging to a long-run limit. When players are somewhat impatient, a complete knowledge transfer is not even attainable at that limit. [^22] This is true in our baseline model. In the model with retirement, knowledge is always fully transferred upon the expert’s retirement, regardless of the discount factor. If the principal had one-period commitment in our setting, however, the first-best outcome would be achievable: The principal, in period 0, should ask the expert to transfer all knowledge to the novice and commit to paying him $\frac{w(0)-w(1)}{1-\delta}$ in period $1$ upon verifying the transfer.
Knowledge Gift
Discount Factor and Speed of Knowledge Transfer
GR and our paper also yield distinct predictions regarding the effect of the discount factor on the speed of knowledge transfer. In GR, greater patience (i.e., a higher discount factor) leads to longer apprenticeships and slower knowledge transfer. This is because an increase in the discount factor raises the marginal value of knowledge to the novice, making him willing to work in exchange for smaller increments of knowledge per period. The expert exploits this by deliberately slowing the transfer of knowledge, thereby extending the apprentice phase during which he captures the full output. To the contrary, in our model, greater patience enhances the principal’s value from future knowledge transfers; as a result, the principal is willing to pay more to the expert in each period, which in turn supports larger and faster knowledge transfers. As discussed in Section 4.1, our prediction seems to align well with the comparison between the long-run cooperation levels observed in the SAIC-GM joint venture and in the SAIC-Volkswagen joint venture.